OpenStudy (anonymous):

HELP PLZZZZ!!!! @iGreen @BloomLocke367 @rafrasan @sleepyjess @TheSmartOne

OpenStudy (anonymous):

What is the exact volume of a cylinder that has a height of 4.9 yd and a radius of 2.2 yd? [ ? ] yd3

OpenStudy (igreen):

\(V = r^2 h\) This is the volume in terms of pi

OpenStudy (igreen):

Plug in 2.2 for 'r' and 4.9 for 'h' and simplify

OpenStudy (anonymous):

10.78?

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@igreen are u there?

OpenStudy (anonymous):

@YesThisIsDog

OpenStudy (anonymous):

oh Hi!

OpenStudy (sleepyjess):

@horsegirl325 , do you still need help?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so i got 10.78

OpenStudy (igreen):

That's incorrect I think

OpenStudy (sleepyjess):

Okay, what do you get after 4.9*2.2?

OpenStudy (igreen):

Yeah..you multiplied 4.9 * 2.2..multiply 2.2 again.

OpenStudy (sleepyjess):

Whoops, I forgot about the square >.<

OpenStudy (anonymous):

sorry forgot ^2. 23.716

OpenStudy (sleepyjess):

Yes, and they want you to use 3.14 for \(\pi\), so * 3.14

OpenStudy (anonymous):

is 23.716 the final answer?

OpenStudy (sleepyjess):

No, multiply by 3.14

OpenStudy (igreen):

Yes, it is @horsegirl325

OpenStudy (igreen):

@sleepyjess It says EXACT

OpenStudy (anonymous):

i did @sleepyjess

OpenStudy (sleepyjess):

No it isn't, they want the exact volume, not in terms of pi

OpenStudy (igreen):

Actually \(23.716 \pi\)

OpenStudy (igreen):

Yes, they do..otherwise they wouldn't say 'exact volume' 'exact volume' means 'in terms of pi'

OpenStudy (igreen):

@horsegirl325 Is there a pi symbol next to the empty box?

OpenStudy (anonymous):

yes

OpenStudy (sleepyjess):

Ok, then I'm wrong

OpenStudy (anonymous):

sorry, it didn't show up when i pasted it:(

OpenStudy (anonymous):

What is the volume of a cone that has a radius of 4 cm and a height of 9 cm? Use 3.14 to approximate pi and express your answer in hundredths. cm3

OpenStudy (anonymous):

what about that 1 @iGreen @sleepyjess

OpenStudy (sleepyjess):

Open a new question :)

OpenStudy (igreen):

1 question per post.

OpenStudy (anonymous):

ok